Two students sit on a seesaw, one student, on the right side of the seesaw, is at 150cm away from the pivot and exerts a force of 320N. The other student is on the opposite side at 120cm and exerts a force of 400N. which side is closer to the ground?
1
Expert's answer
2020-12-14T12:14:55-0500

That side is closer to ground, which experiences bigger torque. By definition, torque is:


"\tau = FL"

where "F" is the force, and "L" is the distance from the pivot. For the right side:


"\tau_1 = 320N\cdot 1.5m = 480\space N\cdot m"

For the left side:


"\tau_2 = 400N\cdot 1.2m = 480\space N\cdot m"

Since "\tau_1 = \tau_2", both sides are at the same level (if both students sat on the seesaw simultaneously).


Answer. Same level.